OpenStudy (anonymous):

Can anyone explain me how to solve this? \[\frac{ d }{ dx }\int\limits_{2\Pi}^{\sqrt[3]{x}} \sin(t)dt=\]

OpenStudy (anonymous):

Apply the fundamental theorem of calculus: \[\large \frac{d}{dx}\int_c^{g(x)}f(t)~dt=f(g(x))\cdot g'(x)\] To see why this works (at least for this example), you can simply compute the integral, then differentiate: \[\Large\begin{align*}\int_{2\pi}^\sqrt[3]x\sin t~dt&=-\left[\cos t\right]_{2\pi}^{\sqrt[3]x}\\ &=-\cos\sqrt[3]x-(-\cos2\pi)\\ &=-\cos\sqrt[3]x+1 \end{align*}\] Differentiating gives \[\large\begin{align*}\frac{d}{dx}[\cdots]&=-\sin\sqrt[3]x\cdot\frac{d}{dx}[\sqrt[3]x]\\ &=\frac{1}{3}\sin\sqrt[3]x \cdot\frac{1}{x^{2/3}} \end{align*}\]

OpenStudy (anonymous):

Notice that in this case, \(g(x)=\sqrt[3]x\) and \(f(t)=\sin t\), which indeed gives \[\large f(g(x))=\sin\sqrt[3]x\] and \(g'(x)=\dfrac{1}{3x^{2/3}}\).

OpenStudy (anonymous):

wait what happens to the +1? before differentiating.

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

What's the derivative of a constant?

OpenStudy (anonymous):

got it. I was anti deriving